OpenStudy (anonymous):

1. 4 and 2/10 x 2 and 2/10 2. 3 and 1/4 x 5 and 1/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wutt??

OpenStudy (anonymous):

The question's not super clear. I think what you mean is: \[\Large\begin{array}{l} 4\frac{2}{{10}} \times 2\frac{2}{{10}}\\ = \frac{{(4 \times 10) + 2}}{{10}} \times \frac{{(2 \times 10) + 2}}{{10}}\\ = \frac{{42}}{{10}} \times \frac{{22}}{{10}}\\ = \frac{{42 \times 22}}{{10 \times 10}}\\ = \frac{{{\rm{924}}}}{{100}} \end{array}\] Try the second one.